• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 08:05:02 on June18,2005

    Entry number 146430

    keyword=Owl Shift Summary

    Crew: V. Sulkosky, L. Lagamba

    00:00 Begin shift with 60 uA on H20 fall target.
    Energy spread < 3E-5
    01:45 Noticed right septum tripped off at 01:20. Reset interlock
    and ramping up.
    02:01 Right septum tripped off again. Asked for beam off.
    Upper coil temp was around 8.1 Kelvin.
    02:30 Right septum ramped up. Upper coil T = 7.2 Kelvin.
    03:40 Right septum temp climbing toward 8 Kelvin. Asked MCC to
    drop current to 55 uA, but we are getting 56 uA. Septum temperature
    has started to drop again.
    05:16 MCC called. They had a box supply trip off, and beam will be
    gone for a while.
    06:19 MCC called and are ready to deliver beam. Asked for 60 uA.
    07:22 Run start failed due to problem with ROC12. Reboot ROC12 and
    ROC16 and restart CODA. Run 4661 is junk!


    Run Summary

    Start-Stop Run Number Events
    ----------------------------------
    23:55-00:28 4651 1M
    00:29-01:02 4652 1M
    01:03-01:49 4653 0.54M right septum tripped
    02:32-03:03 4654 1M
    03:04-03:35 4655 1M
    03:36-04:11 4656 1M current changed to 56 uA during run
    04:14-04:48 4657 1M
    04:49-05:17 4658 0.77M no beam due to box supply problem
    06:21-06:51 4659 1M
    06:52-07:22 4660 1M
    07:28-07:57 4662 1M